If the perimeter of a rectangle is 80ft, what is its area? How would I solve this?
Answer by Alan3354(69443)

If the perimeter of a rectangle is 80ft, what is its area? How would I solve this?
-----------
2W + 2L = 80
W + L = 40
-------------
Area = L*W
There are an infinite # of combinations, each giving a different area.
No unique solution.